a question paper has 8 questions with 4 options per question.in how many ways can one or more of those questions be attempted?

Answer by richard1234(7193)

There are five ways to answer each question (one each for the options, and to leave the question blank).

Hence there are 5*5*...*5 = 5^8 ways to answer the questions. However we have to subtract 1 to account for not attempting any question. Ans = 5^8 - 1